# Practitioner Motives to Use Different Hyperparameter Optimization Methods

### v4

@@ Abstract.
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: TOCHIjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing HCI theory, concepts and modelsccs: Human-centered computing User models
@@ 1. Introduction
HPO by manual tuning is often cumbersome, tedious, and error-prone. To help practitioners with HPO, research with a technical focus (e.g., (Feurer and Hutter, 2019; Bischl et al., 2023)) developed several programmatic HPO methods, including grid search, random search, Bayesian optimization, and evolutionary algorithms, and implemented such methods as software tools (e.g., Hyperopt and Hyperband; (Bergstra et al., 2015; Li et al., 2018a)). Existing HPO methods, programmatic and non-programmatic ones, differ considerably in the way they optimize hyperparameter configurations. Such differences lead some HPO methods to superiority over others. That superiority is usually demonstrated using conventional performance metrics from computer science, including minimization of generalization errors and increase of sample efficiency (Zöller and Huber, 2021; Gijsbers et al., 2022; Lindauer et al., 2022). However, practitioners often use HPO methods that are inferior according to conventional performance metrics (Bouthillier and Varoquaux, 2020). For example, practitioners often prefer to perform grid search over the more sample-efficient Bayesian optimization (Snoek et al., 2012).
@@ 2.1. Hyperparameter Optimization Methods
Manual tuning refers to a set of HPO methods where practitioners choose hyperparameter configurations based on explicit and implicit knowledge and influences of contextual factors. The dependency of manual tuning on practitioners’ experiences and even unconscious heuristics in decisions (Gigerenzer and Brighton, 2009) makes manual tuning very individual to practitioners, rendering the explication of manual tuning difficult. Thus, manual tuning is usually hardly replicable (Musgrave et al., 2020; Simon et al., 2023). Commonly, only intermediate information (e.g.,  tuned hyperparameters) is available, while reasons for selecting hyperparameter configurations often remain unclear. This complicates the formalization of specific HPO methods in manual tuning, making it unclear how practitioners actually proceed. Common strategies include starting optimization from well-performing hyperparameter values (Xin et al., 2021), using a sound experiment design for incremental improvement (Godbole et al., 2023a), and removing irrelevant hyperparameters from the search space (Radosavovic et al., 2020). In addition, difficulty in explication make it hard to evaluate the sample efficiency of strategies in manual tuning. While evidence that manual tuning outperforms advanced HPO methods is lacking, prior publications offer initial evidence that advanced HPO methods can outperform manual tuning in several use cases (Feurer et al., 2016; Melis et al., 2018; Chen et al., 2018; Zhang et al., 2021; Kadra et al., 2021; Zöller et al., 2023).
To tackle shortcomings of manual tuning, AutoML (Feurer and Hutter, 2019) is envisioned to automate all aspects related to development of ML models in a problem-agnostic manner, for example, by programmatic HPO methods. Research on AutoML has primarily approached HPO from a technical perspective (e.g., (Feurer and Hutter, 2019; Bischl et al., 2023)). Typical works (e.g., (Bergstra et al., 2011; Jamieson and Talwalkar, 2016; Golovin et al., 2017; Falkner et al., 2018; Li et al., 2018a; He et al., 2018; Wang et al., 2021c)) focus on performance optimization of ML models in terms of smaller generalization errors, smaller ML model size, or lower latency. Often, such works investigate HPO from a mathematical perspective and treat HPO as a black-box optimization problem: given a problem instance in the form of a dataset and a loss function, a black-box optimizer (e.g., Bayesian optimization) searches for hyperparameter configurations in a predefined search space to enhance an ML model in terms of a given metric (e.g., accuracy on a validation set). Multi-objective optimization methods can be used to specify additional properties of the resulting ML model, such as algorithmic fairness, fast inference, and low model complexity (Gardner et al., 2019; Binder et al., 2020; Karl et al., 2023; Dooley et al., 2022).
@@ 2.2. Research on Hyperparameter Optimization
Extant research on HPO methods is mainly driven by technological advancements that aim at increasing sample efficiency of HPO methods (Bergstra et al., 2011; Hutter et al., 2011; Snoek et al., 2012), reducing time for evaluating objective functions (Swersky et al., 2014; Domhan et al., 2015; Li et al., 2018a), and transferring knowledge from prior optimization runs to similar problem instances (Dyrmishi et al., 2019; Vanschoren, 2019). Such technological advancements are valuable but designed to reach better results in terms of conventional performance metrics. Practitioner motives to use HPO methods beyond conventional performance metrics are largely neglected.
@@
Studies exploring practitioners’ experiences with programmatic HPO methods provide valuable insights into the perceived advantages and disadvantages of these methods (Gil et al., 2019; Wang et al., 2019b, 2021b). Practitioners acknowledge advantages of programmatic HPO methods, which are commonly related to faster turn-around time for building ML models and, thus, higher productivity in developing ML models (Wang et al., 2019b, 2021a). Automatically tuned ML models are often used by practitioners to create initial baseline ML models for subsequent manual tuning or to gain data insights (Wang et al., 2019b, 2021a; Crisan and Fiore-Gartland, 2021). However, AutoML practitioners often bemoan insufficient confidence in the results of programmatic HPO methods and, therefore, refuse to blindly use ML models optimized with programmatic HPO methods (Lee and Macke, 2020; Wang et al., 2019b; Drozdal et al., 2020; Khuat et al., 2023). Even though acknowledging that programmatic HPO is useful to develop well-performing ML models (Wang et al., 2021a), practitioners often refuse to use those HPO methods to not be accountable for ML models they do not understand (Drozdal et al., 2020). A lack of confidence is often linked to the perceived black-box nature of programmatic HPO methods, such as Bayesian optimization and evolutionary algorithms, limiting transparency of optimizer internals. Practitioners prefer support that augments their daily data science work (e.g., through guidance) rather than fully automating it (Crisan and Fiore-Gartland, 2021).
@@
Research on human-guided HPO focuses on involving humans in programmatic HPO methods to improve HPO with dormant domain expertise (Wang et al., 2019b; Higuchi et al., 2021; Wang et al., 2019a). This requires identifying how and when to involve humans in HPO to achieve the best combination of automation and human knowledge (Crisan and Fiore-Gartland, 2021). Especially involvement of practitioners in ML model development, including HPO, seems promising for a higher level of automation. For other tasks, including data acquisition and requirement analysis, practitioners prefer strong human involvement with a low level of automation (Wang et al., 2021b). Interactions of practitioners with software tools for programmatic HPO were structured into different modes of cooperation between practitioners and software tools, ranging from manual tuning to full automation, in the literature (Lee and Macke, 2020; Crisan and Fiore-Gartland, 2021; Wang et al., 2021b).
@@
Extant research describes valuable concepts how practitioners could interact with software tools for programmatic HPO (Sun et al., 2023; Xin et al., 2021) and how to design visual analytics tools for HPO to support practitioners (Higuchi et al., 2021; Zöller et al., 2023; Wang et al., 2019a). Yet, the different programmatic HPO methods are not further differentiated, and practitioner motives to select different HPO methods remain unclear. Supporting better understanding of practitioner motives for selecting HPO methods is the main goal of this study.
@@ 3.1. Semi-structured Interviews with Machine Learning Experts
Academia (14)14(14)( 14 ) Bachelor (2)2(2)( 2 ) <2absent2<2< 2 (4)4(4)( 4 ) ML Innovator (10) CV (8)8(8)( 8 )
Industry (6)6(6)( 6 ) Master (16)16(16)( 16 ) 22224444 (6)6(6)( 6 ) ML Engineer (10) NLP (6)6(6)( 6 )
PhD (2)2(2)( 2 ) 55557777 (7)7(7)( 7 ) RL (5)5(5)( 5 )
>7absent7>7> 7 (3)3(3)( 3 ) TSF (3)3(3)( 3 )
@@ 3.2. Online Survey
Academia (27)27(27)( 27 ) High School (1)1(1)( 1 ) <2absent2<2< 2 (8)8(8)( 8 ) ML Innovator (22) CV (12)12(12)( 12 )
Industry (22)22(22)( 22 ) Bachelor (2)2(2)( 2 ) 22224444 (20)20(20)( 20 ) ML Engineer (19) TD (19)19(19)( 19 )
Master (27)27(27)( 27 ) 55557777 (14)14(14)( 14 ) Novice (8) NLP (15)15(15)( 15 )
Diploma (2)2(2)( 2 ) 888810101010 (4)4(4)( 4 ) TSF (8)8(8)( 8 )
PhD (17)17(17)( 17 ) >10absent10>10> 10 (3)3(3)( 3 ) RL (5)5(5)( 5 )
@@ 5.1. Principal Findings
Besides improving ML model performance in the first place, practitioners are second most interested in decreasing the required amount of necessary computations and their personal efforts to perform HPO, which reflects the basic motivation for development of HPO tools (Wang et al., 2021b). To decrease necessary computations, the study participants predominantly used Bayesian optimization. The frequent use of Bayesian optimization to decrease necessary computations suggests that practitioner perceptions of the benefits of Bayesian optimization are coherent with its benefits empirically shown in prior research (Turner et al., 2021).
@@
To help practitioners increase ML model understanding, various software tools were designed. Such tools, predominantly HPO tools in AutoML, mainly focus on supporting measurements of influences of hyperparameter configurations on ML model performance (e.g., (Hutter et al., 2014; Biedenkapp et al., 2017; Moosbauer et al., 2021; Segel et al., 2023)). With a focus on HPO methods, especially tools for visual analytics are envisioned to support better understanding of internal behaviors of HPO methods by visualizations (e.g., (Golovin et al., 2017; Biedenkapp et al., 2018; Sass et al., 2022; Zöller et al., 2023)). Despite the existence of such tools, practitioners tend to prefer to use manual tuning, which may have different reasons. The first reason may be that practitioners are unaware of HPO tools that can help increase ML model understanding. A second reason may be that HPO tools do not fulfill the information needs of practitioners to increase ML model understanding because HPO tools mainly focus on performance of ML models, which is, as shown in this study, only one practitioner motive in HPO. A third reason may be that the functioning of HPO tools themselves is hard to comprehend for practitioners (e.g.,  because such HPO tools implement unfamiliar and complex HPO methods), which leads practitioners to prefer HPO methods they are familiar with.
@@
The inspection of frequent HPO method and goal combinations revealed that increasing ML model comprehension is the only goal with a strongly higher association with manual tuning than programmatic HPO methods (see Figure 3). A potential explanation could be that established HPO tools usually focus on finding the best performing hyperparameters and do not provide details of other interesting hyperparameter values tested (Xin et al., 2021; Zöller et al., 2023). To overcome this limitation, HPO tools should generate reports on the behavior of different ML models, for example, the importance of individual hyperparameters. For the generation of such reports, many methods are already available, such as functional ANOVA (Hutter et al., 2014), ablation (Biedenkapp et al., 2017), importance of local parameters (Biedenkapp et al., 2018), partial dependence plots (Moosbauer et al., 2021), and symbolic regressions (Segel et al., 2023). Alternatively, HPO tools could provide additional insights about ML model behavior. Especially for more complex search spaces used for building complete ML pipelines, information about the transformation of input data can additionally help increase ML model understanding (Zöller et al., 2023). Including such reports in HPO tools can facilitate leveraging the benefits of advanced HPO methods (e.g.,  high sample efficiency) while still helping practitioners to increase ML model understanding.
@@
Figure 6c shows that ML model comprehension is an important contextual factor for preferring manual tuning instead of a programmatic HPO method. Similarly, multiple interviewees mentioned that they prefer manual tuning when they are confident in predicting the impact of hyperparameters on the ML model behavior. To harness this knowledge, HPO tools should enable practitioners to incorporate their relevant knowledge of behaviors of ML models into HPO tools prior to HPO on a case-by-case basis, tailored to goals like improve ML model performance. For example, practitioners could specify their perceived hyperparameter importance and influences between hyperparameters. Furthermore, ML models comprehension of practitioner could be directly incorporated into the search strategy of HPO methods, for example to reduce search spaces (Xin et al., 2021) and better use available computing resources (the most important contextual factor for Bayesian optimization in Figure 6a). Promising work in this direction includes methods for integrating prior knowledge into Bayesian optimization. This can be achieved by directly specifying priors about the location of the optimum (Li et al., 2020; Ramachandran et al., 2020; Souza et al., 2021; Hvarfner et al., 2022), or structural priors, for example, in the form of log-transformations of hyperparameters (Hutter et al., 2011), monotonicity constraints (Li et al., 2018b), or hyperparameter warping (Snoek et al., 2014).
@@ 5.3. Limitations and Future Work
Future investigations of human decision-making in HPO are a promising research direction that can help to improve AutoML by incorporating human knowledge (Higuchi et al., 2021; Wang et al., 2019a; Sun et al., 2023). The interviewed practitioners reported actions they applied in HPO, which are agnostic to the selected HPO methods, such as choosing a promising subset of hyperparameters to tune and defining corresponding search ranges. Most interviewees proceeded very much alike in choosing HPO methods, selecting hyperparameters, and tuning hyperparameter configurations. Because those interviewees stated to have successfully attained their goals, such similar proceeding HPO allows for the assumption that best practices for actions taken during HPO exist. As the interviewees mainly stated that they largely unconsciously compared HPO methods but achieved satisfactory outcomes, the identification of cognitive heuristics use in practitioners’ decision-making (Gigerenzer and Brighton, 2009; Gigerenzer and Selten, 2002) in HPO appears to be of great potential to advance automation of HPO tasks in AutoML. By identifying functional heuristics that practitioners use in HPO, a better understanding of how practitioners use HPO methods can be reached. This could help enhance AutoML, especially HPO tools, in terms of automating even complex tasks that are still left to human beings by mimicking human decision-making in a resource-efficient manner. Moreover, such heuristics could be helpful to make HPO tools more efficient.

### v5

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: 10.1145/3745771journal: TOCHIjournalvolume: 32journalnumber: 6article: 59publicationmonth: 8ccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing HCI theory, concepts and modelsccs: Human-centered computing User models
@@ 1. Introduction
HPO by manual tuning is often cumbersome, tedious, and error-prone. To help practitioners with HPO, research with a technical focus (e.g., (Feurer and Hutter, 2019; Bischl et al., 2023)) developed several programmatic HPO methods, including grid search, random search, Bayesian optimization, and evolutionary algorithms, and implemented such methods as software tools (e.g., Hyperopt and Hyperband; (Bergstra et al., 2015; Li et al., 2018b)). Existing HPO methods, programmatic and non-programmatic ones, differ considerably in the way they optimize hyperparameter configurations. Such differences lead some HPO methods to superiority over others. That superiority is usually demonstrated using conventional performance metrics from computer science, including minimization of generalization errors and increase of sample efficiency (Zöller and Huber, 2021; Gijsbers et al., 2022; Lindauer et al., 2022). However, practitioners often use HPO methods that are inferior according to conventional performance metrics (Bouthillier and Varoquaux, 2020). For example, practitioners often prefer to perform grid search over the more sample-efficient Bayesian optimization (Snoek et al., 2012).
@@ 2.1. Hyperparameter Optimization Methods
Manual tuning refers to a set of HPO methods where practitioners choose hyperparameter configurations based on explicit and implicit knowledge and influences of contextual factors. The dependency of manual tuning on practitioners’ experiences and even unconscious heuristics in decisions (Gigerenzer and Brighton, 2009) makes manual tuning very individual to practitioners, rendering the explication of manual tuning difficult. Thus, manual tuning is usually hardly replicable (Musgrave et al., 2020; Simon et al., 2023). Commonly, only intermediate information (e.g.,  tuned hyperparameters) is available, while reasons for selecting hyperparameter configurations often remain unclear. This complicates the formalization of specific HPO methods in manual tuning, making it unclear how practitioners actually proceed. Common strategies include starting optimization from well-performing hyperparameter values (Xin et al., 2021), using a sound experiment design for incremental improvement (Godbole et al., 2023b), and removing irrelevant hyperparameters from the search space (Radosavovic et al., 2020). In addition, difficulty in explication make it hard to evaluate the sample efficiency of strategies in manual tuning. While evidence that manual tuning outperforms advanced HPO methods is lacking, prior publications offer initial evidence that advanced HPO methods can outperform manual tuning in several use cases (Feurer et al., 2016; Melis et al., 2018; Chen et al., 2018; Zhang et al., 2021; Kadra et al., 2021; Zöller et al., 2023a).
To tackle shortcomings of manual tuning, AutoML (Feurer and Hutter, 2019) is envisioned to automate all aspects related to development of ML models in a problem-agnostic manner, for example, by programmatic HPO methods. Research on AutoML has primarily approached HPO from a technical perspective (e.g., (Feurer and Hutter, 2019; Bischl et al., 2023)). Typical works (e.g., (Bergstra et al., 2011; Jamieson and Talwalkar, 2016; Golovin et al., 2017; Falkner et al., 2018; Li et al., 2018b; He et al., 2018; Wang et al., 2021a)) focus on performance optimization of ML models in terms of smaller generalization errors, smaller ML model size, or lower latency. Often, such works investigate HPO from a mathematical perspective and treat HPO as a black-box optimization problem: given a problem instance in the form of a dataset and a loss function, a black-box optimizer (e.g., Bayesian optimization) searches for hyperparameter configurations in a predefined search space to enhance an ML model in terms of a given metric (e.g., accuracy on a validation set). Multi-objective optimization methods can be used to specify additional properties of the resulting ML model, such as algorithmic fairness, fast inference, and low model complexity (Gardner et al., 2019; Binder et al., 2020; Karl et al., 2023; Dooley et al., 2022).
@@ 2.2. Research on Hyperparameter Optimization
Extant research on HPO methods is mainly driven by technological advancements that aim at increasing sample efficiency of HPO methods (Bergstra et al., 2011; Hutter et al., 2011; Snoek et al., 2012), reducing time for evaluating objective functions (Swersky et al., 2014; Domhan et al., 2015; Li et al., 2018b), and transferring knowledge from prior optimization runs to similar problem instances (Dyrmishi et al., 2019; Vanschoren, 2019). Such technological advancements are valuable but designed to reach better results in terms of conventional performance metrics. Practitioner motives to use HPO methods beyond conventional performance metrics are largely neglected.
@@
Studies exploring practitioners’ experiences with programmatic HPO methods provide valuable insights into the perceived advantages and disadvantages of these methods (Gil et al., 2019; Wang et al., 2019a, 2021c). Practitioners acknowledge advantages of programmatic HPO methods, which are commonly related to faster turn-around time for building ML models and, thus, higher productivity in developing ML models (Wang et al., 2019a, 2021b). Automatically tuned ML models are often used by practitioners to create initial baseline ML models for subsequent manual tuning or to gain data insights (Wang et al., 2019a, 2021b; Crisan and Fiore-Gartland, 2021). However, AutoML practitioners often bemoan insufficient confidence in the results of programmatic HPO methods and, therefore, refuse to blindly use ML models optimized with programmatic HPO methods (Lee and Macke, 2020; Wang et al., 2019a; Drozdal et al., 2020; Khuat et al., 2023). Even though acknowledging that programmatic HPO is useful to develop well-performing ML models (Wang et al., 2021b), practitioners often refuse to use those HPO methods to not be accountable for ML models they do not understand (Drozdal et al., 2020). A lack of confidence is often linked to the perceived black-box nature of programmatic HPO methods, such as Bayesian optimization and evolutionary algorithms, limiting transparency of optimizer internals. Practitioners prefer support that augments their daily data science work (e.g., through guidance) rather than fully automating it (Crisan and Fiore-Gartland, 2021).
@@
Research on human-guided HPO focuses on involving humans in programmatic HPO methods to improve HPO with dormant domain expertise (Wang et al., 2019a; Higuchi et al., 2021; Wang et al., 2019b). This requires identifying how and when to involve humans in HPO to achieve the best combination of automation and human knowledge (Crisan and Fiore-Gartland, 2021). Especially involvement of practitioners in ML model development, including HPO, seems promising for a higher level of automation. For other tasks, including data acquisition and requirement analysis, practitioners prefer strong human involvement with a low level of automation (Wang et al., 2021c). Interactions of practitioners with software tools for programmatic HPO were structured into different modes of cooperation between practitioners and software tools, ranging from manual tuning to full automation, in the literature (Lee and Macke, 2020; Crisan and Fiore-Gartland, 2021; Wang et al., 2021c).
@@
Extant research describes valuable concepts how practitioners could interact with software tools for programmatic HPO (Sun et al., 2023; Xin et al., 2021) and how to design visual analytics tools for HPO to support practitioners (Higuchi et al., 2021; Zöller et al., 2023b; Wang et al., 2019b). Yet, the different programmatic HPO methods are not further differentiated, and practitioner motives to select different HPO methods remain unclear. Supporting better understanding of practitioner motives for selecting HPO methods is the main goal of this study.
@@ 3.1. Semi-structured Interviews with Machine Learning Experts
Academia $(14)$ Bachelor $(2)$ <2<2 $(4)$ ML Innovator (10) CV $(8)$
Industry $(6)$ Master $(16)$ 2244 $(6)$ ML Engineer (10) NLP $(6)$
PhD $(2)$ 5577 $(7)$ RL $(5)$
>7>7 $(3)$ TSF $(3)$
@@ 3.2. Online Survey
Academia $(27)$ High School $(1)$ <2<2 $(8)$ ML Innovator (22) CV $(12)$
Industry $(22)$ Bachelor $(2)$ 2244 $(20)$ ML Engineer (19) TD $(19)$
Master $(27)$ 5577 $(14)$ Novice (8) NLP $(15)$
Diploma $(2)$ 881010 $(4)$ TSF $(8)$
PhD $(17)$ >10>10 $(3)$ RL $(5)$
@@ 5.1. Principal Findings
Besides improving ML model performance in the first place, practitioners are second most interested in decreasing the required amount of necessary computations and their personal efforts to perform HPO, which reflects the basic motivation for development of HPO tools (Wang et al., 2021c). To decrease necessary computations, the study participants predominantly used Bayesian optimization. The frequent use of Bayesian optimization to decrease necessary computations suggests that practitioner perceptions of the benefits of Bayesian optimization are coherent with its benefits empirically shown in prior research (Turner et al., 2021).
@@
To help practitioners increase ML model understanding, various software tools were designed. Such tools, predominantly HPO tools in AutoML, mainly focus on supporting measurements of influences of hyperparameter configurations on ML model performance (e.g., (Hutter et al., 2014; Biedenkapp et al., 2017; Moosbauer et al., 2021; Segel et al., 2023)). With a focus on HPO methods, especially tools for visual analytics are envisioned to support better understanding of internal behaviors of HPO methods by visualizations (e.g., (Golovin et al., 2017; Biedenkapp et al., 2018; Sass et al., 2022; Zöller et al., 2023b)). Despite the existence of such tools, practitioners tend to prefer to use manual tuning, which may have different reasons. The first reason may be that practitioners are unaware of HPO tools that can help increase ML model understanding. A second reason may be that HPO tools do not fulfill the information needs of practitioners to increase ML model understanding because HPO tools mainly focus on performance of ML models, which is, as shown in this study, only one practitioner motive in HPO. A third reason may be that the functioning of HPO tools themselves is hard to comprehend for practitioners (e.g.,  because such HPO tools implement unfamiliar and complex HPO methods), which leads practitioners to prefer HPO methods they are familiar with.
@@
The inspection of frequent HPO method and goal combinations revealed that increasing ML model comprehension is the only goal with a strongly higher association with manual tuning than programmatic HPO methods (see Figure 3). A potential explanation could be that established HPO tools usually focus on finding the best performing hyperparameters and do not provide details of other interesting hyperparameter values tested (Xin et al., 2021; Zöller et al., 2023b). To overcome this limitation, HPO tools should generate reports on the behavior of different ML models, for example, the importance of individual hyperparameters. For the generation of such reports, many methods are already available, such as functional ANOVA (Hutter et al., 2014), ablation (Biedenkapp et al., 2017), importance of local parameters (Biedenkapp et al., 2018), partial dependence plots (Moosbauer et al., 2021), and symbolic regressions (Segel et al., 2023). Alternatively, HPO tools could provide additional insights about ML model behavior. Especially for more complex search spaces used for building complete ML pipelines, information about the transformation of input data can additionally help increase ML model understanding (Zöller et al., 2023b). Including such reports in HPO tools can facilitate leveraging the benefits of advanced HPO methods (e.g.,  high sample efficiency) while still helping practitioners to increase ML model understanding.
@@
Figure 6c shows that ML model comprehension is an important contextual factor for preferring manual tuning instead of a programmatic HPO method. Similarly, multiple interviewees mentioned that they prefer manual tuning when they are confident in predicting the impact of hyperparameters on the ML model behavior. To harness this knowledge, HPO tools should enable practitioners to incorporate their relevant knowledge of behaviors of ML models into HPO tools prior to HPO on a case-by-case basis, tailored to goals like improve ML model performance. For example, practitioners could specify their perceived hyperparameter importance and influences between hyperparameters. Furthermore, ML models comprehension of practitioner could be directly incorporated into the search strategy of HPO methods, for example to reduce search spaces (Xin et al., 2021) and better use available computing resources (the most important contextual factor for Bayesian optimization in Figure 6a). Promising work in this direction includes methods for integrating prior knowledge into Bayesian optimization. This can be achieved by directly specifying priors about the location of the optimum (Li et al., 2020; Ramachandran et al., 2020; Souza et al., 2021; Hvarfner et al., 2022), or structural priors, for example, in the form of log-transformations of hyperparameters (Hutter et al., 2011), monotonicity constraints (Li et al., 2018a), or hyperparameter warping (Snoek et al., 2014).
@@ 5.3. Limitations and Future Work
Future investigations of human decision-making in HPO are a promising research direction that can help to improve AutoML by incorporating human knowledge (Higuchi et al., 2021; Wang et al., 2019b; Sun et al., 2023). The interviewed practitioners reported actions they applied in HPO, which are agnostic to the selected HPO methods, such as choosing a promising subset of hyperparameters to tune and defining corresponding search ranges. Most interviewees proceeded very much alike in choosing HPO methods, selecting hyperparameters, and tuning hyperparameter configurations. Because those interviewees stated to have successfully attained their goals, such similar proceeding HPO allows for the assumption that best practices for actions taken during HPO exist. As the interviewees mainly stated that they largely unconsciously compared HPO methods but achieved satisfactory outcomes, the identification of cognitive heuristics use in practitioners’ decision-making (Gigerenzer and Brighton, 2009; Gigerenzer and Selten, 2002) in HPO appears to be of great potential to advance automation of HPO tasks in AutoML. By identifying functional heuristics that practitioners use in HPO, a better understanding of how practitioners use HPO methods can be reached. This could help enhance AutoML, especially HPO tools, in terms of automating even complex tasks that are still left to human beings by mimicking human decision-making in a resource-efficient manner. Moreover, such heuristics could be helpful to make HPO tools more efficient.